Experiment 13 Acid-Base Equilibria, Part I1 Lab partner Report Sheet Data Titration 1: Solution being titrated is Q500 ml of 100 MHA. mI, were required to reach the phenolphthalein endpoint Titration 2: Solution being titraed is 25ml ml of MHA Volume of base p pH Volume of base 0.00 mL 05.0% 5.52 ISoom 592 L.96 Construct a graph with pH on the y-axis and volume of base added on the x-axs. The electronic endpoint is the mid-point of the steepest part of the curve. This occurred when exactly mL base had been added. Use your graph to obtain this value. This should be clearly marked on your graph. Use a dotted line from the electronic endpoint to the x-axis to show this. See question #5

media%2F6ad%2F6adb4d5b-b445-4cc5-944e-f5
Answer #1

Plotting a titration curve withe the values of pH and volume of the base:

Titration curve 14 12 10 8 6 4 0 10 15 20 25 Vbase (mL) → Endpoint volume= 15.9mL

The electronic end point occurs exactly when 15.9 mL of base has been added.

1. The exact concentration of the base by using phenolphtalein, can be calculated using:

M_A*V_A=M_B*V_B

Solving for MB:

M_A*(V_A/V_B)=M_B

Where MA is 0.100 M, VA is 25 mL and VB is 16.01 mL, the exact concentration of base is:

0.100M*(25mL/16.01mL)=M_B= 0.156M

2. The exact concentration of the base by electronic end point, can be calculated using:

M_A*V_A=M_B*V_B

Solving for MB:

M_A*(V_A/V_B)=M_B

Where MA is 0.100 M, VA is 25 mL and VB is 15.9 mL, the exact concentration of base is:

0.100M*(25mL/15.9mL)=M_B= 0.157M

3. The percent difference between these values, using (2) as the standard is calculated using:

%E = \frac{0.157-0.156}{0.157}*100 = 0.64%

This shows us that the use of indicators in acid-base titrations may cause a little error due to its character as an acid-base substance, and the pH useful range of it.

4. The useful range for phenolphtalein is 8.3-10 pH units. It is right in the region when electronic end point is reached, so we can think that the error in titration is low, as we calculate in 3.

It was the expected because in the plot of titration curve, the pH changes near to the end point are between 6 an 12 pH units, and the indicator has an useful range of 8.3-10. Also, end point in the titration curve is closely near to 8.3 pH units.